OpenStudy (anonymous):

When solving systems of 3 equations & 3 unknowns..... If they all cancel does that mean is No Solution?

OpenStudy (unklerhaukus):

for example x=x y=y z=z there are three fold infinite solutions

OpenStudy (anonymous):

Ok. I'm confused

OpenStudy (unklerhaukus):

no solutions occur when the equations dont make sense x=1 y=2 x+y=0 has zero solutions

OpenStudy (anonymous):

the problem is: 4x +2y+6z=13, -12x+3y-5z=8, -4x+7y+7z=34. When I put two equations together to eliminate 1 variable. My results end up the same for Eq4 & 5 which then cancel each other. Ex. -9y-13z=-47 9y+13z=47

OpenStudy (anonymous):

Oh. Ok so since it all cancels. Then the answer is infinite solution?

OpenStudy (unklerhaukus):

4x +2y+6z=13, i -12x+3y-5z=8, ii -4x+7y+7z=34, iii 9y+13z=47 i+iii

OpenStudy (anonymous):

I got what you were trying to explain. It's infinite solution because 0=0 which mean it is dependent

OpenStudy (unklerhaukus):

how did you get to 0=0?

OpenStudy (anonymous):

The all cancel out so variables are 0 and = 0

OpenStudy (anonymous):

I did it on my paper. I had to multiply Eq 4 by -1 to utilize elimination method and 4 and 5 cancel out

OpenStudy (unklerhaukus):

4x +2y+6z=13, i -12x+3y-5z=8, ii 9y+13z=47 ii+3i

OpenStudy (unklerhaukus):

i+iii - ii+3i 0x+0y+0z=0 all values of x,y,z satisfy this equation

OpenStudy (anonymous):

so the answer is dependent

OpenStudy (anonymous):

infinite solutions

OpenStudy (unklerhaukus):

if you graph the three plains , you will find that they look something like this  the intersection is a infinite line of solutions